OpenStudy (anonymous):

9^lnx-4*3^lnx+1 + 27=0

OpenStudy (zehanz):

Remember 9=3², so the formula can be rewritten as: \(3^{2^{\ln x}}-4\cdot3^{\ln x}+27=0\) and then, because \(3^{2^{\ln x}}=3^{2{\ln x}}=(3^{\ln x})^2\): \((3^{\ln x})^2-4\cdot3^{\ln x}+27=0\). Doesn't that remind you of quadratic equations (set \(p=3^{\ln x}\))?

OpenStudy (zehanz):

I used the rule: \((a^{b})^c=a^{bc}=a^{cb}=(a^{c})^b\)

OpenStudy (anonymous):

3^2lnx−4⋅3^lnx+1+27=0 ...you did not put +1 in force ..is the same thing withaout it??

OpenStudy (zehanz):

Not quite :( I overlooked that 1. Should it be ...^ln(x+1)+27 =... ? Otherwise it would just be ...^ln(x) + 28 = ...

OpenStudy (zehanz):

Oh, I see now what it has to be: \((3^{\ln (x)})^2-4 \cdot 3^{\ln(x) +1}+27=0\), so this is \((3^{\ln (x)})^2-4 \cdot 3^{\ln(x)}\cdot 3^1+27=0\) \((3^{\ln (x)})^2-12 \cdot 3^{\ln(x)}+27=0\) Now you can factor it into \((3^{\ln x}-9)\cdot(3^{\ln x}-3)=0\), which gives \(3^{\ln x}=9\) or \(3^{\ln x}=3\), meaning \(\ln x=2\) or \(\ln x=1\), so \(x=e^2\) or \(x=e\)